OpenStudy (anonymous):

Find slope and Y intercept of a line whose equation is 12x+4y+8 need step by step explanation please :)

OpenStudy (anonymous):

shouldn't it be 12x + 4y + 8 = 0?

OpenStudy (anonymous):

i assume it is set to 0. it didn't say though... im in math 65 so its usually not complicated

OpenStudy (anonymous):

Hi there, if you're trying to find a y-inercept, you have to put a zero where the x is. This is so you can isolate y. Same for the x intercept. (Just vice versa.)

OpenStudy (anonymous):

so is y simply 2?

OpenStudy (anonymous):

Well ok let's see: First we should isolate (y) so it looks like an equation for a line: 12x + 4y + 8 = 0 4y + 8 = -12x 4y = -12x - 8 y = -3x - 2 the y-intercept is just the y-value when x=0, so let's go ahead and put in x=0 y = -3(0) - 2 y = -2 so the intercept is -2 the slope is just the number in front of (x), in this case that is -3 so for y=-3x - 2 y-int: -2 slope (or m): -3

OpenStudy (anonymous):

why is y -2 and not positive?

OpenStudy (anonymous):

You are subtracting the 2 from the rest of the equation. It would be positive if you were adding the 2.

OpenStudy (anonymous):

Thank you it makes sense :)

OpenStudy (anonymous):

because when x = 0, y = -2, here's how the graph would look like